I need the solution in 20 minutes please (urgent!)

I have a test in 20 minutes and I just need to know this problem's solution beforehand as I was just told a similar one will be on it for sure.
Use th uation Method to write the equations for shear and moment for each beam segment for the beam below. Draw the FBD for each of the beam segments. Clearly label the distance x on the FBD. You do not need to draw the diagrams. 800 Nm 500 N O B I m le 0,5m 0,5 m
Answer #1

To draw a free body diagram for each of segments which can be shown below as :

800 N/m Ax X1

Using the equation method, the equations for shear and moment for each beam segment which is given as :

\sum MA = 0 ;    (800 N/m) (1 m) (0.5 m) + (500 N) (1.5 m) - By (2 m) = 0

By (2 m) = [(400 N.m) + (750 N.m)]

By = (1450 N.m) / (2 m)

By = 725 N

\sum Fy = 0 ;    Ay - (800 N/m) (1 m) - (500 N) + (725 N) = 0

Ay = [(800 N) + (500 N) - (725 N)]

Ay = 575 N

\sum Fx = 0 ;   Ax = 0 N